The sum of john's age and Jorge's age is 56. Ten years ago John was twice as old as Jorge. How old is Jorge now?
Answer by stanbon(75887)

The sum of john's age and Jorge's age is 56. Ten years ago John was twice as old as Jorge. How old is Jorge now?
-----------------------------
Equations:
h + g = 56
(h-10) = 2(g-10)
------------".
-------------
Solve for "h and for "g"
----
Cheers,
Stan H.